OpenStudy (anonymous):

A 0.26-kg rock is thrown vertically upward from the top of a cliff that is 30 m high. When it hits the ground at the base of th e cliff, the rock has a speed of 33 m/s . Assuming that air resistance can be ignored, find the initial speed of the rock.

OpenStudy (lgbasallote):

you can split it up into two parts: 1) going up 2) going down going up: initial speed = v1 final speed = 0 acceleration = -9.8 distance = d going down: initial speed = 0 final speed = 33 acceleration = 9.8 distance = d + 30 so the equation for going up is \[V_2 ^ 2 = V_1 ^ 2 + 2as\] substitute.. \[0 = V_1^2 - 19.6d\] \[19.6d = V_1^2\] \[d = \frac{V_1^2}{19.6}\] now equation for going down \[V_2 ^2 = V_1^2 + 2as\] \[33^2 = 0 +9.8(d+30)\] \[1089 = 9.8(\frac{V_1^2}{19.6}) + 294\] \[795 =9.8(\frac{V_1^2}{19.6})\] \[81.12 = \frac{V_1^2}{19.6}\] \[1590 = V_1^2\] \[39.87 = V_1\]

OpenStudy (lgbasallote):

did you follow that?

OpenStudy (anonymous):

yeah , thanks a lot . you sure are a help Igbasallote

OpenStudy (lgbasallote):

welcome